# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 4 Fast Ground Completion Introducing Additional Constants
For each symbol $\sigma\in\Sigma\langle E,F\rangle$, we pad the binary form of its number with leading zeros to length $\ell$. The binary sequence obtained in this way is called the binary representation of the symbol $\sigma\in\Sigma\langle E,F\rangle$, and is denoted as $\phi(\sigma)$. Similarly, for each constant $a\in C\langle E,F\rangle$, we pad the binary form of its number with leading zeros to length $\ell$. The binary sequence obtained in this way is called the binary representation of the constant $a\in C\langle E,F\rangle$, and is denoted as $\phi(a)$.
The binary representation of a rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E,F\rangle$ is the string $\phi(\sigma)\phi(a_{1}),\ldots,\phi(a_{m})\phi(a)$. We store the binary representations of the rewrite rules in $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT1$ in lexicographic order. The binary representation of an ordered pair $(a,b)\in BSTEP\langle E,F\rangle$ is the ordered pair $(\phi(a),\phi(b))$. We store the binary representations of the elements of $BSTEP\langle E,F\rangle$ without duplicates in the red-black tree $RBT2$ in lexicographic order. Here for every ordered pair $(a,b)\in BSTEP\langle E,F\rangle$, the search key is the first component $\phi(a)$, the second component, $\phi(b)$ is satellite data.
var $i$, $m$: ${\mathbb{N}}$; $\sigma$: $\Sigma$; $a,a_{1},\ldots,a_{mar_{\Sigma}}:C\langle E,F\rangle$; $RBT1$, $RBT2$: red-black tree;
@@
$a:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x)$;
@@
$a_{i}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x/i)$;
$R\langle E,F\rangle:=R\langle E,F\rangle\cup\{\,\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\,\}$;
$BSTEP\langle E,F\rangle:=BSTEP\langle E,F\rangle\cup\{\,(a_{1},a),\ldots,(a_{m},a)\,\}$;
@@
By Statement 4.45 and Proposition 2.2, the time complexity of storing $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT1$ is $O(n\,\mathrm{log}\,n)$, where $n=size(E)+size(F)$. By Statement 4.45 and Proposition 2.2, the time complexity of storing $BSTEP\langle E,F\rangle$ without duplicates in the red-black tree $RBT2$ is $O(n\,\mathrm{log}\,n)$, where $n=size(E)+size(F)$.
Proposition 4.47.
[2] Algorithm FGC takes as input a signature $\Sigma$, and GTESs $E$ and $F$ over $\Sigma$. Let $n=size(E)+size(F)$.
@@
2. In $O(n\,\mathrm{log}\,n)$ time, Algorithm FGC computes the GTRS $R\langle E,F\rangle$ over $\Sigma\cup C\langle E,F\rangle$, and stores $R\langle E,F\rangle$ without duplicates in the lexical order $<_{lex}$ on binary sequences in the red-black tree $RBT1$.
@@
Fülöp and Vágvölgyi [2] showed that we can construct the GTRS $R\langle E,F\rangle$ in $O(n\,\mathrm{log}\;n)$ time. Hence we can construct the relation $BSTEP\langle E,F\rangle$ in $O(n\,\mathrm{log}\;n)$ time as well. In addition to the result of Fülöp and Vágvölgyi [2], we now compute the cost of storing the GTRS $R\langle E,F\rangle$. and the cost of storing the relation $BSTEP\langle E,F\rangle$. First we show Statement 1. Algorithm FGC produces the set of nullary symbols $C\langle E,F\rangle$. By Definition 4.26, for every equivalence class $\vartheta\in\Theta\langle E,F\rangle$, Algorithm FGC gives a unique name for $\vartheta$ which is the unique name of the $\rho\langle E,F\rangle$ equivalence class $\eta$ such that $\vartheta=\{\,\widehat{x}\mid x\in\eta\,\}$.
We now show Statement 2. Algorithm FGC produces the GTRS $R\langle E,F\rangle$ apart from that, for every equivalence class $\vartheta\in\Theta\langle E,F\rangle$, Algorithm FGC writes the unique name for $\vartheta$ [2]. By Statement 4.45, $size(R\langle E,F\rangle)\leq 2\cdot|W|\leq 2\cdot n$. Then by Proposition 2.2, the time-complexity of inserting the elements of $R\langle E\cup F,\emptyset\rangle$ into an empty red-black tree, not allowing duplicate values in the resulting red-black tree $RBT1$, is $O(n\,\mathrm{log}\;n)$,
We now show Statement 3. In the relation $BSTEP\langle E,F\rangle\subseteq C\langle E,F\rangle\times C\langle E,F\rangle$, we collect all pairs $(a_{i},a)$ such that there is a rewrite rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a$ in $R\langle E,F\rangle$. By Statement 6.63, $|BSTEP\langle E,F\rangle|\leq n$. Then by Proposition 2.2, the time-complexity of inserting the elements of $BSTEP\langle E,F\rangle$ into an empty red-black tree, not allowing duplicate values in the resulting red-black tree, is $O(n\,\mathrm{log}\;n)$.
@@
Let $E$ and $F$ be GTESs over a ranked alphabet $\Sigma$.
@@
If $FLAG\langle E,F\rangle=\mathrm{false}$, then return $\mathrm{false}$. If $R\langle E,F\rangle=\emptyset$, then return $\mathrm{false}$. Otherwise, traverse the red-black tree $RBT$ in in-order traversal, and read all binary sequences stored in the red-black tree $RBT$ in the lexical order $<_{lex}$ on binary sequences. For every $m\in{\mathbb{N}}$, $\sigma\in\Sigma\langle R\langle E,F\rangle_{m}$, read the rules with $\sigma$ appering on the left-hand side one after another. During this process,
@@
count all rules with $\sigma$ appearing on the left-hand side, and store the result in the variable $count$;
@@
Algorithm DTOT decides in $O(n)$ time whether $R\langle E,F\rangle$ is total, where $n=size(E)+size(F)$.
@@
First, assume that $R\langle E,F\rangle=\emptyset$. Recall that there exists a constant $\sigma\in\Sigma_{0}$. Then there does not exist a rewrite rule with left-hand side $\sigma$ in $R\langle E,F\rangle$. Consequently, $R\langle E,F\rangle$ is not total.
@@
Second, assume that $R\langle E,F\rangle\neq\emptyset$. If $FLAG\langle E,F\rangle=\mathrm{false}$, that is, $|\Sigma\langle E,F\rangle|<|\Sigma|$, then $R\langle E,F\rangle$ is not total. Assume that $FLAG\langle E,F\rangle=\mathrm{true}$, that is, $|\Sigma\langle E,F\rangle|=|\Sigma|$. Then for each symbol $\sigma\in\Sigma$, we count the rewrite rules in $R\langle E,F\rangle$ with $\sigma$ appearing in the left-hand side and store the result in the variable $count$. If there exists a symbol $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, such that for $\sigma$, $count<|\langle C\langle E,F\rangle\rangle|^{m}$, then there exist constants $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$ such that there does not exist a rewrite rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R\langle E,F\rangle$. Therefore, $R\langle E,F\rangle$ is not total. Conversely if $R\langle E,F\rangle$ is not total, then there exists a symbol $\sigma\in\Sigma_{m}$ and constants $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$ such that there does not exist a rewrite rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R\langle E,F\rangle$. Then for $\sigma$, $count<|\langle C\langle E,F\rangle\rangle|^{m}$. Consequently if $R\langle E,F\rangle$ is total, then Algorithm DTOT outputs true; otherwise it returns false.
@@
We now consider the time complexity of the Algorithm DTOT. If $|\Sigma\langle E,F\rangle|<|\Sigma|$ or $R\langle E,F\rangle=\emptyset$, then Algorithm DTOT takes less than $O(n)$ time. Assume that $|\Sigma\langle E,F\rangle|=|\Sigma|$ and $R\langle E,F\rangle\neq\emptyset$. Then Algorithm DTOT traverse the red-black tree $RBT$ in in-order traversal, and it reads all binary sequences stored in the red-black tree $RBT$ in the lexical order $<_{lex}$ on binary sequences. For every $\sigma\in\Sigma\langle R\langle E,F\rangle$, it reads the rules with $\sigma$ appering on the left-hand side one after another, and counts the rewrite rules in $R\langle E,F\rangle$ with $\sigma$ appearing in the left-hand side for all symbols $\sigma\in\Sigma$. This takes $O(n)$ time.
@@
If there exists a symbol $\sigma\in\Sigma$ such that $\sigma$ does not appear in $R\langle E,F\rangle$, then Algorithm DTOT outpouts false, and does not evaluate the expression $count<|C\langle E,F\rangle|^{m}$. Otherwise, for every symbol $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, we carry out $m$ multiplications when computing $|C\langle E,F\rangle|^{m}$. Because each symbol $\sigma\in\Sigma\langle R\langle E,F\rangle_{m}$, $m\in{\mathbb{N}}$, has $m$ constants in its arguments in the left-hand side of some rewrite rule in $R\langle E,F\rangle$.
The number of multiplications $\leq$ $\sum_{m\in{\mathbb{N}}}\sum_{\sigma\in\Sigma\langle R\langle E,F\rangle_{m}}m\leq$ $size(R\langle E,F\rangle)\leq n$.
by Statement 4.45. Consequently, the number of multiplications is less than or equal to $n$. The number of comparisons of nonnegative integers is $|\Sigma|\leq n$. Therefore the time complexity of algorithm DTOT is $O(n)$. ∎
@@ 5 Running Examples Continued
We now illustrate the concepts of Section 4 by continuing our running examples. As in Section 3, throughout this section $E$ and $F$ are GTESs over a signature $\Sigma$, and $W$ stands for $ST\langle E\cup F\rangle$.
@@
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E,F\rangle},\;\newline g(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;g(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow},\;g(\$)/_{\Theta\langle E,F\rangle}\;\}$,
@@
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E,F\rangle},\;\newline g(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;g(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E,F\rangle}\;\}$,
@@
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline g(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},g(\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
@@
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
We note that $R\langle E,F\rangle$, $R\langle F,E\rangle$, and $R\langle E\cup F,E\rangle$ are total.

### v2

@@ 4 Fast Ground Completion Introducing Additional Constants
For each symbol $\sigma\in\Sigma\langle E,F\rangle$, we pad the binary form of its number with leading zeros to length $\ell$. The binary sequence obtained in this way is called the binary representation of the symbol $\sigma\in\Sigma\langle E,F\rangle$, and is denoted as $\phi(\sigma)$. Similarly, for each constant ${b}\in C\langle E,F\rangle$, we pad the binary form of its number with leading zeros to length $\ell$. The binary sequence obtained in this way is called the binary representation of the constant ${b}\in C\langle E,F\rangle$, and is denoted as $\phi({b})$.
The binary representation of a rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E,F\rangle$ is the string $\phi(\sigma)\phi({b}_{1}),\ldots,\phi({b}_{m})\phi({b})$. We store the binary representations of the rewrite rules in $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT1$ in lexicographic order. The binary representation of an ordered pair $({b},{c})\in BSTEP\langle E,F\rangle$ is the ordered pair $(\phi({b}),\phi({c}))$. We store the binary representations of the elements of $BSTEP\langle E,F\rangle$ without duplicates in the red-black tree $RBT2$ in lexicographic order. Here for every ordered pair $({b},{c})\in BSTEP\langle E,F\rangle$, the search key is the first component $\phi({b})$, the second component, $\phi({c})$ is satellite data.
var $i$, $m$: $\mathbb{N}$}; $\sigma$: $\Sigma$; ${b},{b}_{1},\ldots,{b}_{mar_{\Sigma}}:C\langle E,F\rangle$; $RBT1$, $RBT2$: red-black tree;
@@
${b}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x)$;
@@
${b}_{i}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x/i)$;
$R\langle E,F\rangle:=R\langle E,F\rangle\cup\{\,\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\,\}$;
$BSTEP\langle E,F\rangle:=BSTEP\langle E,F\rangle\cup\{\,({b}_{1},{b}),\ldots,({b}_{m},{b})\,\}$;
@@
By Statement 4.46 and Proposition 2.2, the time complexity of storing $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT1$ is $O(n\,\mathrm{log}\,n)$, and the time complexity of storing $BSTEP\langle E,F\rangle$ without duplicates in the red-black tree $RBT2$ is $O(n\,\mathrm{log}\,n)$, where $n=\mathrm{size}(E)+\mathrm{size}(F)$.
Proposition 4.48 ([5, Algorithm 3.6]).
Algorithm FGC takes as input a signature $\Sigma$, and GTESs $E$ and $F$ over $\Sigma$. Let $n=\mathrm{size}(E)+\mathrm{size}(F)$.
@@
2. In $O(n\,\mathrm{log}\,n)$ time, Algorithm FGC computes the GTRS $R\langle E,F\rangle$ over $\Sigma\cup C\langle E,F\rangle$, and stores $R\langle E,F\rangle$ without duplicates in the lexicographic order $_{lex}$ on binary sequences in the red-black tree $RBT1$.
@@
Fülöp and Vágvölgyi [5] showed that we can construct the GTRS $R\langle E,F\rangle$ in $O(n\,\mathrm{log}\,n)$ time. Hence we can construct the relation $BSTEP\langle E,F\rangle$ in $O(n\,\mathrm{log}\,n)$ time as well. In addition to the result of Fülöp and Vágvölgyi [5], we now compute the cost of storing the GTRS $R\langle E,F\rangle$, and the cost of storing the relation $BSTEP\langle E,F\rangle$. First we show Statement 1. Algorithm FGC produces the set of nullary symbols $C\langle E,F\rangle$. By Definition 4.27, for every equivalence class $\vartheta\in\Theta\langle E,F\rangle$, Algorithm FGC gives a unique name for $\vartheta$ which is the unique name of the $\rho\langle E,F\rangle$ equivalence class $\eta$ such that $\vartheta=\{\,\widehat{x}\mid x\in\eta\,\}$.
We now show Statement 2. Algorithm FGC produces the GTRS $R\langle E,F\rangle$ apart from that, for every equivalence class $\vartheta\in\Theta\langle E,F\rangle$, Algorithm FGC writes the unique name for $\vartheta$ [5]. By Statement 4.46, $\mathrm{size}(R\langle E,F\rangle)\leq 2\cdot|W|\leq 2n$. Then by Proposition 2.2, the time-complexity of inserting the elements of $R\langle E\cup F,\emptyset\rangle$ into an empty red-black tree, not allowing duplicate values in the resulting red-black tree $RBT1$, is $O(n\,\mathrm{log}\,n)$,
We now show Statement 3. In the relation $BSTEP\langle E,F\rangle\subseteq C\langle E,F\rangle\times C\langle E,F\rangle$, we collect all pairs $({b}_{i},{b})$ such that there is a rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$ in $R\langle E,F\rangle$. By Statement 4.46, $|BSTEP\langle E,F\rangle|\leq 2n$. Then by Proposition 2.2, the time-complexity of inserting the elements of $BSTEP\langle E,F\rangle$ into an empty red-black tree, not allowing duplicate values in the resulting red-black tree, is $O(n\,\mathrm{log}\,n)$. ∎
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$.
@@
If $FLAG\langle E,F\rangle=\mathrm{false}$, then return $\mathrm{false}$. If $R\langle E,F\rangle=\emptyset$, then return $\mathrm{false}$. Otherwise, traverse the red-black tree $RBT$ in in-order traversal, and read all binary sequences stored in the red-black tree $RBT$ in the lexicographic order $_{lex}$ on binary sequences. For every $m\in{\mathbb{N}}$, $\sigma\in\Sigma\langle R\langle E,F\rangle_{m}$, read the rules with $\sigma$ appering on the left-hand side one after another. During this process,
@@
count all rules with $\sigma$ appearing on the left-hand side, and store the result in the variable countcount;
@@
Algorithm DTOT decides in $O(n)$ time whether $R\langle E,F\rangle$ is total, where $n=\mathrm{size}(E)+\mathrm{size}(F)$.
@@
If $R\langle E,F\rangle=\emptyset$, then by Statement 4.40, we are done.
@@
Assume that $R\langle E,F\rangle\neq\emptyset$. If $FLAG\langle E,F\rangle=\mathrm{false}$, that is, $|\Sigma\langle E,F\rangle|<|\Sigma|$, then $R\langle E,F\rangle$ is not total. Assume that $FLAG\langle E,F\rangle=\mathrm{true}$, that is, $|\Sigma\langle E,F\rangle|=|\Sigma|$. Then for each symbol $\sigma\in\Sigma$, we count the rewrite rules in $R\langle E,F\rangle$ with $\sigma$ appearing in the left-hand side and store the result in the variable countcount. If there exists a symbol $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, such that for $\sigma$, $count<|\langle C\langle E,F\rangle\rangle|^{m}$, then there exist constants ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$ such that there does not exist a rewrite rule with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$ in $R\langle E,F\rangle$. Therefore, $R\langle E,F\rangle$ is not total. Conversely if $R\langle E,F\rangle$ is not total, then there exists a symbol $\sigma\in\Sigma_{m}$ and constants ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$ such that there does not exist a rewrite rule with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$ in $R\langle E,F\rangle$. Then for $\sigma$, $count<|\langle C\langle E,F\rangle\rangle|^{m}$. Consequently if $R\langle E,F\rangle$ is total, then Algorithm DTOT outputs true; otherwise it returns false.
@@
We now consider the time complexity of the Algorithm DTOT. If $|\Sigma\langle E,F\rangle|<|\Sigma|$ or $R\langle E,F\rangle=\emptyset$, then Algorithm DTOT takes less than $O(n)$ time. Assume that $|\Sigma\langle E,F\rangle|=|\Sigma|$ and $R\langle E,F\rangle\neq\emptyset$. Then Algorithm DTOT traverse the red-black tree $RBT$ in in-order traversal, and it reads all binary sequences stored in the red-black tree $RBT$ in the lexicographic order $_{lex}$ on binary sequences. For every $\sigma\in\Sigma\langle R\langle E,F\rangle\rangle$, it reads the rules with $\sigma$ appering on the left-hand side one after another, and counts the rewrite rules in $R\langle E,F\rangle$ with $\sigma$ appearing in the left-hand side for all symbols $\sigma\in\Sigma$. This takes $O(n)$ time.
@@
If there exists a symbol $\sigma\in\Sigma$ such that $\sigma$ does not appear in $R\langle E,F\rangle$, then Algorithm DTOT outpouts false, and does not evaluate the expression $count<|C\langle E,F\rangle|^{m}$. Otherwise, for every symbol $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, we carry out $m$ multiplications when computing $|C\langle E,F\rangle|^{m}$. Because each symbol $\sigma\in\Sigma\langle R\langle E,F\rangle\rangle_{m}$, $m\in{\mathbb{N}}$, has $m$ constants in its arguments in the left-hand side of some rewrite rule in $R\langle E,F\rangle$. The number of multiplications is less than or equal to $\sum_{m\in{\mathbb{N}}}\sum_{\sigma\in\Sigma\langle R\langle E,F\rangle\rangle_{m}}m\leq$ $\mathrm{size}(R\langle E,F\rangle)\leq 2n$ by Statement 4.46. Consequently, the number of multiplications is less than or equal to $2n$. The number of comparisons of nonnegative integers is $|\Sigma|\leq n$. Therefore the time complexity of algorithm DTOT is $O(n)$. ∎
@@ 5 Running Examples Continued
We now illustrate the concepts of Section 4 by continuing our running examples. We compute the sets of constants $C\langle E,F\rangle$, $C\langle F,E\rangle$, $C\langle E\cup F,\emptyset\rangle$, and the GTRSs $R\langle E,F\rangle$, $R\langle F,E\rangle$, $R\langle E\cup F,\emptyset\rangle$.
Recall that $E$ and $F$ are GTESs over a signature $\Sigma$, and $W$ stands for $ST\langle E\cup F\rangle$.
@@
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;f(\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\mathdollar)/_{\Theta\langle E,F\rangle},\;\newline g(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;g(\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow},\;g(\mathdollar)/_{\Theta\langle E,F\rangle}\;\}$,
@@
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle F,E\rangle},\;f(\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar)/_{\Theta\langle F,E\rangle},\;\newline g(\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle F,E\rangle},\;g(\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar)/_{\Theta\langle F,E\rangle}\;\}$,
@@
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline g(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},g(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
@@
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.